# A Survey on Future Frame Synthesis: Bridging Deterministic and Generative Approaches

### v4

@@ 6.1 Diffusion-based generation
Diffusion models [47] have become dominant in the image generation realm. The Latent Diffusion Model (LDM) [81] explores the generation capability in the latent space of images, significantly improving computational efficiency and reducing costs. LDM has paved the way for introducing diffusion models to the video generation realm. The extensions of LDM exhibit robust generative capabilities [100]. Furthermore, Video LDM [9] leverages pre-trained image models for video generation, offering multi-modal, high-resolution, long-term video generation ability. Similarly, SEINE [20] introduces a versatile video diffusion model, enabling the creation of transition sequences to generate longer videos from shorter clips.
@@
Recently, there has been a surge in research efforts aimed at harnessing additional information in conjunction with RGB images to fulfill the text-guided video completion task. LFDM [75] builds upon the latent diffusion models to synthesize optical flow sequences in latent space based on textual guidance. Seer [40] inflates Stable Diffusion [81] along the temporal axis, enabling the model to utilize natural language instructions and reference frames to imagine multiple variations of future outcomes. Emu Video [37] takes a different approach by initially generating an image conditioned on textual guidance and subsequently extrapolating it into a video. Consequently, this method can be flexibly applied to create video based on a given image with different textual inputs. DynamiCrafter [119] broadens the application of text-guided image animation to open-domain images. SparseCtrl [42] allows for sketch-to-video generation, depth-to-video generation, and video prediction with an expanded input range.
@@
Some works attempt to provide a more precise understanding of the guidance of text and preserve the guidance information along the time dimension. MicroCinema [107] adopts a divide-and-conquer strategy to tackle appearance and temporal coherence challenges. This model employs a two-stage generation pipeline, which initially utilizes any existing text-to-image generator for the creation of the initial image, and subsequently introduces a dedicated text-guided video generation framework specifically designed for motion modeling. LivePhoto [19] introduces a framework that incorporates motion intensity as a supplementary factor to enhance control over desired motions. Additionally, it proposes a text re-weighting mechanism to emphasize motion descriptions, showcasing impressive performance on text-guided video synthesis tasks. I2VGen-XL [131] utilizes static images for semantic and qualitative guidance, showcasing the diverse approaches in text-guided video synthesis research.
@@
Although diffusion models have almost become the most popular generative models, many researchers are attempting to preserve the architecture of autoregressive models to simultaneously utilize the advantages of both mechanisms [113]. Despite the limitations in data availability and model scalability that hindered the early generative video synthesis algorithms mentioned above, they laid the groundwork for integrating LLMs and diffusion models in the field of video generation. Recently, GAIA-1 [49] and Sora [11] leverage the strengths of both diffusion models and large language models to achieve more creative, general, and scalable video synthesis.
@@ 6.2 Token-based generation
In the field of image/video generation, while diffusion-based methods attract the largest community and most researchers, these algorithms generally have much smaller model sizes compared to the large-scale language models (LLMs) commonly seen today. As a result, many studies focus on exploring how transformers can accomplish these tasks. The advantage of this approach lies in leveraging a large amount of related research and optimization methods from LLMs to validate the scaling laws in the visual domain. Implementing transformers for future frame synthesis involves two key components: an efficiently scalable LLM framework, and a good image/video tokenizer. Innovations like VQ-VAE [96] and VQGAN [30] combine autoregressive models and adversarial training strategies to tackle image quantization and tokenization. However, an effective visual tokenizer should minimize the number of tokens per image or video clip segment while ensuring near-lossless visual reconstruction. During LLM training, the typical context size ranges from 8k to 32k tokens, while lossless reconstruction of a $256\times 256$ resolution image usually requires no fewer than 500 tokens with extremely large vocabulary size [127], with high-resolution images demanding thousands of tokens. This makes it difficult for the model to process long sequences of video clips during training, severely limiting the model’s video generation capabilities.
@@
Before the rise of LLMs [2], the transformative impact of transformers has been evident in time-series modeling. Video Transformer [112] applies transformers to create an autoregressive model, pioneering the application of transformer architectures in the context of video synthesis. While this autoregressive framework marked a successful attempt in the realm of natural video generation, it also inherited the consistent drawbacks of transformers, namely extensive training resource consumption and prolonged inference time. The Latent Video Transformer (LVT) [78] proposes a novel approach by considering the modeling of dynamics and predicting future features in an autoregressive manner within a latent space to release the computation burden. The NUWA [115] framework proposes a versatile 3D transformer encoder-decoder architecture, adaptable to various data modalities and tasks, further highlighting the potential of transformers in video synthesis. NUWA-Infinity [67] expands on this with an innovative generation mechanism, with the ambitious goal of achieving infinite high-resolution video generation. This progress signifies a continuous effort to unify generative tasks across different modalities, showcasing the evolving landscape of future frame synthesis research.
@@
LVM [5] introduces sequential modeling to expand the learning capabilities of large-scale vision models, showcasing the scalability and flexibility of sequence models in handling in-context learning. The paper also proposes a concept of ”visual sentence” where a sequence of images with intrinsic relationships is arranged similarly to a sentence in language, allowing the model to leverage the information within the sequence to perform sentence continuation. This unified manner enables the model to utilize the prompt information of the visual sentence to synthesize future frames or complete other visual tasks without relying on any non-pixel-level knowledge. The in-context learning is not unfamiliar in the visual domain. Painter [103] introduces a general framework for visual learning that enables images to ”speak” through in-context visual learning, enhancing both image generation and understanding capabilities. Similarly, SegGPT [105] investigates using a GPT architecture for image segmentation and proposes the concept of segmenting everything demonstrating the possibility of achieving unified image segmentation in an unsupervised learning environment, further advancing generalization in visual segmentation tasks.
@@
In the video generation domain, MAGVIT [127] presents a masked generative video Transformer that efficiently processes video by masking certain parts and generating the missing segments. Building on this, MAGVIT-v2 [128] proposes that language models can outperform diffusion models in visual generation tasks, emphasizing the critical role of visual tokenizers in visual generation, which offers new insights into the future development of visual generation tasks. Utilizing the powerful visual tokenizer proposed by MAGVIT-v2, VideoPoet [59] introduces a large language model for zero-shot video generation, pushing the boundaries of unsupervised learning in the video generation field. It allows users to generate or modify videos based on high-level textual prompts. The model excels in capturing the temporal and contextual relationships within video data, enabling it to perform tasks such as video generation, continuation, and editing directly from text descriptions. These studies collectively integrate visual tokenizers and large language models to offer unified and scalable frameworks for visual learning, driving breakthroughs in future frame synthesis tasks.
@@
As mentioned above, text-guided generative video synthesis algorithms are designed to create a sequence of frames by integrating context frames and textual guidance. Text-guided Video Completion (TVC) task involves completing videos based on various conditions, including the first frame (video prediction task), the last frame (video rewind task), or a combination of both (video transition task), all of which are guided by textual instructions. MMVG [34] addresses the TVC task by utilizing autoregressive encoder-decoder architectures, integrating text and frame features, resulting in a unified framework capable of handling multiple video synthesis tasks.
7 Future research
Based on the overall development trends in the fields of Artificial Intelligence and Computer Vision, and considering the challenges in the future frame synthesis domain discussed in Section 2.2, we contend that the trajectory of video synthesis research should diverge from the traditional trends of data simplification and model lightweight in low-level tasks. The imperative should be to harness substantial computational resources and high-resolution, long-duration video data. Future research should prioritize the formulation of evaluation metrics that incentivize stochastic synthesis, thus expanding the potential of simulation of the human world. The ultimate objective is to cultivate video synthesis models with a profound comprehension of the inherent dynamics within videos. Such models could execute video generation over extended time horizons with high stochastic complexity in the real world.
Dataset Year Category # Videos # Clip Frames Resolution Extra Annotations
KTH Action [83] 2004 Human 2, 391 95 160 $\times$ 120 Class
Caltech Pedestrian [26] 2009 Human 137 1, 824 640 $\times$ 480 Bounding Box
HMDB51 [60] 2011 Human 6, 766 93 414 $\times$ 404$\dagger$ Class
UCF101 [89] 2012 Human 13, 320 187 320 $\times$ 240 Class
J-HMDB [54] 2013 Human 928 34 320 $\times$ 240 OF, Ins, HJ, Class
KITTI [36] 2013 Traffic 151 323 1242 $\times$ 375 OF, BBox, Sem, Ins, Depth
Penn Action [132] 2013 Human 2, 326 70 480 $\times$ 270$\dagger$ Human Joint, Class
SJTU 4K [88] 2013 General 15 300 3840 $\times$ 2160 -
Sports-1M [57] 2014 Human 1, 133, 158 variable variable Class
Moving MNIST [90] 2015 Simulation 10, 000 20 64 $\times$ 64 -
Cityscapes [22] 2016 Traffic 46 869 2048 $\times$ 1024 Semantic, Instance, Depth
YouTube-8M [1] 2016 General 8, 200, 000 variable variable Class
Robotic Pushing [31] 2016 Robot 59, 000 25 640 $\times$ 512 Class
DAVIS17 [77] 2017 General 150 73 3840 $\times$ 2026$\dagger$ Semantic
Something-Something [38] 2017 Object 220, 847 45 427 $\times$ 240$\dagger$ Text
ShapeStacks [39] 2018 Simulation 36, 000 16 224 $\times$ 224 Semantic
SM-MNIST [24] 2018 Simulation customize customize 64 $\times$ 64 -
D2-City [17] 2019 Traffic 11, 211 750 1080p / 720p BBox
Kinetics-700 [15, 14] 2019 Human 650, 000 250 variable Class
RoboNet [23] 2019 Robot 161, 000 93 64 $\times$ 48 -
Vimeo-90K [121] 2019 General 91, 701 7 448 $\times$ 256 -
BDD100K [126] 2020 Traffic 100, 000 1175 1280 $\times$ 720 BBox, Semantic, Depth
nuScenes [12] 2020 Traffic 1, 000 40 1600 $\times$ 900 BBox, Semantic
WebVid [6] 2021 General 10, 732, 607 449 596 $\times$ 336 Text
X4K1000FPS [87] 2021 General 4, 408 65 4096 $\times$ 2160 -
SportsSlomo [18] 2024 Human 130,000 7 $1280\times 720$
InternVideo2 [109] 2024 General 2, 000, 000 variable variable Text, Action
OpenDV-YouTube [123] 2024 Traffic 2, 139 100,000+ variable Text
denotes the mean value. $\dagger$ denotes the median value.
@@ 8 Conclusion
In this survey, we have discussed various aspects of future frame synthesis, covering key topics such as predominant datasets, evolving algorithms, current challenges, and future trends. We propose a novel taxonomy based on algorithm stochasticity, showcasing a significant shift from deterministic approaches toward generative methodologies. This survey emphasizes the importance of balancing pixel-level accuracy with a deep understanding of complex scene dynamics in video synthesis. Additionally, we delve into the intricacies of stochastic birth-and-death phenomena, advocating for enhanced evaluation metrics and the utilization of substantial computational resources and large-scale video datasets. These insights aim to guide future research directions in video synthesis. As the field advances, we anticipate the development of models that provide a more profound and nuanced grasp of nature in the real world. This evolution promises to enhance accuracy, efficiency, and creativity, paving the way for novel applications and research.
Acknowledgments
This work is supported by Beijing Natural Science Foundation L231014.
Appendix
9 Datasets
We offer an overview of the most widely used datasets in video synthesis, highlighting their amounts of data and additional supervisory modalities, thereby providing a comprehensive picture of the current dataset landscape in this field.
10 Application realms
The applications of future frame synthesis span a wide range of domains, showcasing its significance in different fields.
World Model.
World models [43, 134] provide a more general framework for simulating and predicting the behavior of complex systems. A world model is often used in reinforcement learning and robotics to enable agents to make informed decisions and take actions that lead to desired outcomes. To construct a world model, future frame synthesis is an important learning objective [44, 45, 106].
Autonomous driving.
Future frame synthesis is vital for self-driving cars and drones to anticipate the movement of objects, pedestrians, and other vehicles, ensuring safe navigation. GAIA-1 [49] employs a unified world model, combining multi-modal large language models and diffusion procedures to predict control signals and future frames.
Robot navigation.
Future frame synthesis is employed in guiding robots through dynamic environments, allowing them to plan their paths, move objects, and avoid obstacles effectively [32].
Cinema industry.
Future frame synthesis is utilized in special effects, animation, and pre-visualization, aiding filmmakers in creating realistic scenes and enhancing the overall cinematic experience [73].
Meteorological community.
Future frame synthesis plays a crucial role in weather forecasting, helping meteorologists simulate and predict atmospheric conditions, contributing to improved accuracy in weather predictions [86].

### v5

@@ 5.1 Diffusion-Based Generation
Diffusion models (Ho et al., 2020) have emerged as the dominant approach in image generation. The Latent Diffusion Model (LDM) (Rombach et al., 2022) extends this capability into the latent space of images, significantly enhancing computational efficiency and reducing costs. This innovation has paved the way for the introduction of diffusion models into the realm of video generation.
Latent diffusion model extensions.
@@
Extensions of LDM have demonstrated robust generative capabilities in video generation (Voleti et al., 2022). For instance, Video LDM (Blattmann et al., 2023b) leverages pre-trained image models to generate videos, offering multi-modal, high-resolution, and long-term video generation capabilities. Similarly, SEINE (Chen et al., 2024a) introduces a versatile video diffusion model that creates transition sequences, enabling the generation of longer videos from shorter clips.
Text-guided video completion with additional information.
@@
Recent research efforts have focused on harnessing additional information along with RGB images to fulfill the text-guided video completion task. LFDM (Ni et al., 2023) extends latent diffusion models to synthesize optical flow sequences in latent space based on textual guidance. Seer (Gu et al., 2023) inflates Stable Diffusion (Rombach et al., 2022) along the temporal axis, allowing the model to use natural language instructions and reference frames to imagine multiple variations of future outcomes. Emu Video (Girdhar et al., 2023) generates an image conditioned on textual guidance and then extrapolates it into a video, making it flexible for creating videos based on different textual inputs. DynamiCrafter (Xing et al., 2023a) broadens the application of text-guided image animation to open-domain images. SparseCtrl (Guo et al., 2023) enables sketch-to-video generation, depth-to-video generation, and video prediction with an expanded input range. There are also methods, such as PEEKABOO Jain et al. (2024), that attempt to achieve interactive synthesis, aiming to open up unprecedented applications and creativity.
Preserving text guidance.
@@
Some works aim to provide a more precise understanding of text guidance and preserve this information along the time dimension. MicroCinema (Wang et al., 2023d) adopts a divide-and-conquer strategy to address appearance and temporal coherence challenges. It uses a two-stage generation pipeline, initially creating the initial image using any existing text-to-image generator and then introducing a dedicated text-guided video generation framework for motion modeling. LivePhoto (Chen et al., 2023) introduces a framework that incorporates motion intensity as a supplementary factor to enhance control over desired motions. It also proposes a text re-weighting mechanism to emphasize motion descriptions, demonstrating impressive performance in text-guided video synthesis tasks. I2VGen-XL (Zhang et al., 2023) utilizes static images for semantic and qualitative guidance, showcasing diverse approaches in text-guided video synthesis research.
Integrating autoregressive models.
@@
Despite diffusion models becoming the most popular generative models, some researchers attempt to preserve the architecture of autoregressive models to simultaneously utilize the advantages of both mechanisms (Weng et al., 2023). Early generative video synthesis algorithms faced limitations in data availability and model scalability, but they laid the groundwork for integrating LLMs and diffusion models in video generation. Recently, GAIA-1 (Hu et al., 2023a) and Sora (Brooks et al., 2024) leverage the strengths of both diffusion models and LLMs to achieve more creative, general, and scalable video synthesis.
These advancements in video generation highlight the potential of diffusion models to create high-quality, controllable, and diverse video content, pushing the boundaries of what is possible in the field of computer vision and artificial intelligence.
Challenges.
Diffusion models have made significant strides in video generation, yet several critical challenges persist. Maintaining temporal coherence and consistency across frames is essential for realism but remains difficult (Chen et al., 2024b; Xu et al., 2024). The computational efficiency and scalability of these models are hindered by their resource-intensive nature, limiting their widespread adoption (Peebles & Xie, 2023). Controllability and interpretability issues arise as textual guidance may not always align with visual outcomes, and model behaviors can be opaque. Data availability and diversity are crucial for training robust models, but obtaining comprehensive datasets is challenging.
@@ 5.2 Token-Based Generation
In the realm of image/video generation, diffusion-based methods have garnered significant attention. However, these algorithms typically possess smaller model sizes compared to contemporary large-scale language models (LLMs). Recent research has also placed significant emphasis on exploring how LLMs can accomplish these tasks, leveraging extensive related research and optimization methods from LLMs to validate scaling laws in the visual domain.
Key components.
Implementing transformers FFS involves two crucial elements: an efficiently scalable LLM framework and a proficient image/video tokenizer. Innovations such as VQ-VAE (Van Den Oord et al., 2017) and VQGAN (Esser et al., 2021) have combined auto-regressive models with adversarial training strategies to address image quantization and tokenization. An effective visual tokenizer should minimize tokens per image or video clip segment while ensuring near-lossless visual reconstruction. However, the high token requirements for lossless reconstruction, especially for high-resolution images, pose challenges for processing long video sequences during training, limiting video generation capabilities.
Early transformer applications.
Before the advent of LLMs (Brown, 2020; Achiam et al., 2023), transformers have already made a significant impact on time-series modeling. Video Transformer (Weissenborn et al., 2019) pioneers the application of transformer architectures in video synthesis, creating an auto-regressive model. Despite its success, it inherits the common drawbacks of transformers: extensive training resource consumption and prolonged inference time.
Latent space modeling.
@@
The Latent Video Transformer (LVT) (Rakhimov et al., 2020) introduces a novel approach by modeling dynamics and predicting future features auto-regressively within a latent space, reducing the computational burden. The NUWA framework (Wu et al., 2022a) proposes a versatile 3D transformer encoder-decoder architecture adaptable to various data modalities and tasks, further showcasing the potential of transformers in video synthesis. NUWA-Infinity (Liang et al., 2022) expands on this with an innovative generation mechanism aimed at achieving infinite high-resolution video generation, reflecting ongoing efforts to unify generative tasks across different modalities.
Sequential modeling and visual sentences.
@@
LVM (Bai et al., 2023) introduces sequential modeling to enhance the learning capabilities of large-scale vision models, demonstrating the scalability and flexibility of sequence models in in-context learning. The concept of "visual sentence" was proposed, where a sequence of images with intrinsic relationships is arranged similarly to a sentence in the language, enabling the model to leverage sequence information for sentence continuation and other visual tasks without relying on non-pixel-level knowledge.
In-context learning in visual domain.
In-context learning is not new to the visual domain. Painter (Wang et al., 2023b) introduces a general framework for visual learning that enables images to "speak" through in-context visual learning, enhancing both image generation and understanding. SegGPT (Wang et al., 2023c) investigates using a GPT architecture for image segmentation, proposing the concept of segmenting everything and demonstrating the possibility of achieving unified image segmentation in an unsupervised learning environment, advancing generalization in visual segmentation tasks.
Advances in video generation.
@@
In the video generation domain, MAGVIT (Yu et al., 2023a) presents a masked generative video Transformer that efficiently processes video by masking certain parts and generating missing segments. MAGVIT-v2 (Yu et al., 2023b) suggests that transformers might outperform diffusion models in visual generation tasks, emphasizing the critical role of visual tokenizers. VideoPoet (Kondratyuk et al., 2023) introduces a large language model for zero-shot video generation, pushing the boundaries of unsupervised learning in video generation. It allows users to generate or modify videos based on high-level textual prompts, excelling in capturing temporal and contextual relationships within video data.
@@
Text-guided generative video synthesis algorithms create a sequence of frames by integrating context frames and textual guidance. The Text-guided Video Completion (TVC) task involves completing videos based on various conditions, including the first frame (video prediction task), the last frame (video rewind task), or a combination of both (video transition task), guided by textual instructions. MMVG (Fu et al., 2023) addresses the TVC task by utilizing auto-regressive encoder-decoder architectures, integrating text and frame features, resulting in a unified framework capable of handling multiple video synthesis tasks.
These studies collectively integrate visual tokenizers and large language models to offer unified and scalable frameworks for visual learning, driving breakthroughs in FFS tasks. The evolving landscape of FFS research continues to showcase the potential of transformers and LLMs in unifying generative tasks across different modalities.
Challenges.
Token-based generation for FFS faces significant hurdles, including efficient visual tokenizer to balance minimal tokens with near-lossless reconstruction, particularly for high-resolution content. The computational resource demands of transformer-based models pose adoption barriers, especially in resource-constrained settings. Because when the computing power budget is insufficient, we may not observe the emergence phenomenon reported in previous studies under this paradigm (Bai et al., 2024). Recently,  Sun et al. (2024a) suggest that one reason token-based models cannot match the visual quality of diffusion models is the lack of integration of high-quality community assets, such as training infrastructure and high-quality data.
6 Application Realms
The applications of FFS span a wide range of domains, showcasing its significance in different fields.
World model.
World models (Ha & Schmidhuber, 2018; Zhu et al., 2024) provide a more general framework for simulating and predicting the behavior of complex systems. A world model is often used in reinforcement learning and robotics to enable agents to make informed decisions and take actions that lead to desired outcomes. To construct a world model, FFS is an important learning objective (Hafner et al., 2020; 2023a; Wang et al., 2024a).  Escontrela et al. (2024) show that video prediction can also serve as a part of reward modeling to aid reinforcement learning.
Autonomous driving.
FFS is indispensable for autonomous vehicles, including self-driving cars and drones, as it allows them to anticipate the movement of objects, pedestrians, and other vehicles. This capability is crucial for ensuring safe navigation. For instance, GAIA-1 (Hu et al., 2023a) leverages a unified world model that integrates multi-modal large language models and diffusion processes to predict control signals and future frames, enhancing the vehicle’s decision-making capabilities. Currently, how to effectively utilize image information to aid trajectory prediction in real-world scenarios remains a challenge (Nayakanti et al., 2023; Varadarajan et al., 2022).
Robot navigation.
In robotics, FFS is employed to guide robots through dynamic environments. It enables robots to plan their paths, move objects, and avoid obstacles effectively, as demonstrated by Finn & Levine (2017). By predicting future states, robots can make proactive decisions, enhancing their adaptability and efficiency in complex settings.
Cinema industry.
FFS finds application in the cinema industry, particularly in special effects, animation, and pre-visualization. It aids filmmakers in creating realistic scenes and enhancing the overall cinematic experience. For example, Mahapatra & Kulkarni (2022) utilize FFS to generate visually compelling sequences that contribute to the storytelling and artistic expression in films.
Meteorological community.
FFS plays a crucial role in weather forecasting, aiding meteorologists in simulating and predicting atmospheric conditions. By accurately predicting future frames, FFS contributes to improved accuracy in weather predictions, as exemplified by the work of Shi et al. (2017). This capability is vital for weather forecasting services and disaster preparedness.
Anomaly detection.
Liu et al. (2018) propose an approach to anomaly detection in videos by leveraging future frame prediction, where normal events are expected to be predictable while abnormal ones are not. The method introduces a motion constraint in addition to appearance constraints to ensure the predicted future frames are consistent with the ground truth in both spatial and temporal aspects.
Overall, these diverse applications underscore the importance and potential of FFS as a tool for understanding and interacting with the world around us. Its ability to predict future states from past observations makes it a valuable asset in fields ranging from artificial intelligence to entertainment and beyond.
Challenges.
When migrating the FFS methods to applications, potential issues that may be encountered. For instance, in specific domains, we might need to explore solutions for few-shot learning (Gui et al., 2018) or adaptation during inference (Choi et al., 2021). We also need to study how to reasonably integrate vectorized data with image data. Moreover, in real-world tasks, people often prefer methods with good interpretability, while most end-to-end FFS methods are not easily interpretable.
7 Related Work
Previous Surveys on Video Prediction.
The field of video prediction and related topics, such as action recognition and spatio-temporal predictive learning, has seen significant advancements in recent years, largely driven by deep learning techniques. Several comprehensive surveys have provided an overview of the state-of-the-art methods, datasets, and evaluation metrics in this domain.  Zhou et al. (2020) introduce next-frame prediction networks before 2020, categorizing them into sequence-to-one and sequence-to-sequence architectures. The paper compares these approaches by analyzing network architecture and loss function design, providing a quantitative performance comparison using off-the-shelf datasets and evaluation metrics.  Oprea et al. (2020) offer a review of deep learning methods for video prediction, defining the fundamentals and background concepts, and analyzing existing models according to a proposed taxonomy. The survey includes experimental results to facilitate the assessment of the state of the art on a quantitative basis. Rasouli (2020a) provide an overview of vision-based prediction algorithms, focusing on deep learning approaches. The paper categorizes these algorithms into video prediction, action prediction, trajectory prediction, body motion prediction, and other prediction applications, discussing common architectures, training methods, data types, evaluation metrics, and datasets.  Kong & Fu (2022) survey the state-of-the-art techniques in action recognition and prediction, discussing existing models, popular algorithms, technical difficulties, action databases, evaluation protocols, and future directions.  Tan et al. (2023) propose OpenSTL, a comprehensive benchmark for spatio-temporal predictive learning, categorizing approaches into recurrent-based and recurrent-free models. The paper conducts standard evaluations on various datasets and provides an analysis of how model architecture and dataset properties affect performance.
Surveys on Video Diffusion Models.
Xing et al. (2023b) present a comprehensive review of video diffusion models in the AIGC era, categorizing the work into video generation, video editing, and other video understanding tasks. The survey includes a thorough review of the literature in these areas and discusses challenges and future trends.  Li et al. (2024) present the survey of recent advancements in long video generation, summarizing them into two key paradigms: divide and conquer, and temporal auto-regressive. It also provides a comprehensive overview and classification of datasets and evaluation metrics, discussing emerging challenges and future directions in this dynamic field.  Sun et al. (2024b) review Sora, OpenAI’s text-to-video model, categorizing literature into evolutionary generators, excellent pursuit, and realistic panorama, while discussing datasets, metrics, challenges, and future research directions.
Our Survey Focus.
Our survey comprehensively reviews historical and contemporary works in FFS, focusing on the transition from deterministic to generative synthesis methodologies. The survey highlights significant advancements and shifts in approach, emphasizing the growing importance of generative models in achieving realistic and diverse future frame predictions.
@@ 8 Conclusion
In this survey, we have discussed various aspects of FFS, covering key topics such as predominant datasets, evolving algorithms and current challenges.
Based on the overall development trends in the field of artificial intelligence, we believe that the trajectory of video synthesis research should diverge. On one hand, research should focus on the trend towards model lightweighting suitable for high-definition video applications, targeting low-level objectives such as video compression and short-term motion estimation. On the other hand, research should explore how models can fundamentally understand the laws of the world and create content by harnessing substantial computational resources and diverse, long-duration video data. For the latter, future research should prioritize the formulation of evaluation metrics that incentivize stochastic synthesis, thus expanding the potential of simulation of the human world. The ultimate objective is to cultivate video synthesis models with a profound comprehension of the inherent dynamics within videos. Such models could execute video generation over extended time horizons with high stochastic complexity in the real world.
Our taxonomy is based on algorithm stochasticity, showcasing a significant shift from deterministic approaches toward generative methodologies. This survey emphasizes the importance of balancing pixel-level accuracy with a deep understanding of complex scene dynamics in video synthesis. Additionally, we delve into the intricacies of stochastic birth-and-death phenomena, advocating for enhanced evaluation metrics and the utilization of substantial computational resources and large-scale video datasets. We also sort out the different research directions under this topic and discuss the current challenges. These insights aim to guide future research directions in video synthesis. As the field advances, we anticipate the development of models that provide a more profound and nuanced grasp of nature in the real world. This evolution promises to enhance accuracy, efficiency, and creativity, paving the way for novel applications and research.